# Blowing star formation away in AGN Hosts (BAH) - I. First Observation of Warm Molecular Outflows with JWST MIRI

### v1

@@ Abstract
We use the James Webb Space Telescope (JWST) Mid-Infrared Instrument (MIRI) medium-resolution spectrometer (MRS) observations of the radio loud AGN host UGC 8782 to map the warm molecular and ionized gas kinematics. The data reveal outflows in the inner 2 kpc, seen in low ionization (traced by the [Ar ii] 6.99 m emission) and in warm molecular gas (traced by the H2 rotational transitions), observed for the first time in an AGN host. We find a maximum mass-outflow rate of M⊙ yr-1 at 900 pc from the nucleus for the warm outflow (198 KT1000 K) and estimate and outflow rate of up to M⊙ yr-1 for the hotter gas phase (T1000 K). These outflows can clear the entire nuclear reservoir of warm molecular gas in about 1 Myr. The derived kinetic power of the molecular outflows lead to coupling efficiencies of 2-5 percent of the AGN luminosity, way above the minimum expected to the AGN feedback be effective quenching the star formation.
@@ 1 Introduction
The molecular gas is the main driver of star formation (SF) (Kennicutt & Evans, 2012) and AGN activity (Storchi-Bergmann & Schnorr-Müller, 2019). Conventionally, it is common to separate molecular gas into three phases with well-defined tracers: (i) The cold gas, in which the temperatures reaches no more then 100 K. This phase is the densest gas phase and is commonly traced by (sub)millimeter transitions of molecules such as CO and HCN; (ii) The warm molecular gas phase, 100 K 1000 K, summing up to 1-30 percent of total molecular gas mass (Roussel et al., 2007). This specific gas phase can be traced by the rotational H2 transitions seen in the mid-infrared (MIR); (iii) The hot molecular phase, K, mainly traced by the roto-vibrational H2 transitions in the near-infrared. This gas phase represents only of the mass of the cold molecular gas phase (Dale et al., 2005; Emonts et al., 2014; Pereira-Santaella et al., 2016). Spatially resolved observations of the structure and kinematics of both cold and hot molecular gas emissions in the central regions of galaxies have been meticulously obtained through ALMA and near-infrared integral field unit observations (Lamperti et al., 2022; Riffel et al., 2023b).
@@ 3 Results and discussions
The emission of the outflow component (bottom panels of Fig. 2) is spatially resolved (with FWHM of its flux distribution of about 2.4 times that of the PSF) and concentrated mostly around the MIR nucleus and radio core. In addition, we simulated a rotating disk convolved with the PSF to assess the impact of beam smearing on the observed line profiles, which was found to be negligible (refer to the Appendix A). This component is blue-shifted relative to the systemic velocity of the galaxy and presents values of up to 600 km s-1. The median velocity and values are 70 km s-1 and 455 km s-1 for the [Ar ii] and 50 km s-1 and 430 km s-1 for the H2, respectively. From the observed H2 emission and kinematics, we derive the warm molecular gas outflow properties, for the first time, in UGC 8782.
We measure the gas mass and temperature using the transitions H2 (0-0) S(1),H2 (0-0) S(2) , H2 (0-0) S(3), H2 (0-0) S(4) and H2 (0-0) S(5), as our data does not include the higher order transitions, as well as the H2 (0-0) S(0) transition. We correct the observed fluxes for extinction using a visual extinction of mag, the mean value for the outflow component (Riffel et al., 2023a) and adopting the extinction law from Gordon et al. (2023).
@@
Given the AGN bolometric luminosity of erg s-1 (Riffel et al., 2023a), the molecular outflow demonstrates maximum kinetic coupling efficiencies () ranging from 2–5% of the AGN luminosity, the kinectic power of the hot and warm phase are respectively erg s-1 and erg s-1. This range far surpasses the minimum required for effective AGN feedback in suppressing star formation (Hopkins & Elvis, 2010), particularly when considering the potential presence of a cold molecular gas reservoir that may also contribute to the outflow. Nevertheless, the outflows in UGC 8782 are likely mechanically driven by the radio jet, as the kinetic power of the outflow is one order of magnitude lower than the object’s radio source power of 1043 erg s-1 (Lanz et al., 2015). Considering a constant outflow rate the warm and hot H2 gas reservoir would be depleted in around Myr.
@@ 4 Conclusions
We reported, for the first time, the observations of warm molecular gas outflows from the AGN of UGC 8782. We find that the mass outflow rate in warm molecular gas is substantially higher than that observed in ionized gas across all radii, with coupling efficiencies exceeding the minimum required for AGN feedback to be effective in quenching star formation. The observed molecular outflows in UGC 8782 of up to 6 M⊙yr-1 are able of redistributing all the warm molecular gas available in the inner 2 kpc of the UGC 8782 in only 1 Myr.

### v2

# Blowing star formation away in AGN Hosts (BAH) - I. Observation of Warm Molecular Outflows with JWST MIRI

@@ Abstract
We use the James Webb Space Telescope (JWST) Mid-Infrared Instrument (MIRI) medium-resolution spectrometer (MRS) observations of the radio loud AGN host UGC 8782 to map the warm molecular and ionized gas kinematics. The data reveal spatially resolved outflows in the inner 2 kpc, seen in low ionization (traced by the [Ar ii] 6.99 m emission) and in warm molecular gas (traced by the H2 rotational transitions). We find a maximum mass-outflow rate of M⊙ yr-1 at 900 pc from the nucleus for the warm outflow (198 KT1000 K) and estimate and outflow rate of up to M⊙ yr-1 for the hotter gas phase (T1000 K). These outflows can clear the entire nuclear reservoir of warm molecular gas in about 1 Myr. The derived kinetic power of the molecular outflows lead to coupling efficiencies of 2-5 percent of the AGN luminosity, way above the minimum expected to the AGN feedback be effective quenching the star formation.
@@ 1 Introduction
The molecular gas is the main driver of star formation (SF) (Kennicutt & Evans, 2012) and AGN activity (Storchi-Bergmann & Schnorr-Müller, 2019). Conventionally, it is common to separate molecular gas into three phases with well-defined tracers: (i) The cold gas, in which the temperatures reaches no more then 100 K. This phase is the densest gas phase and is commonly traced by (sub)millimeter transitions of molecules such as CO and HCN; (ii) The warm molecular gas phase, 100 K 1000 K, summing up to 1-30 percent of total molecular gas mass (Roussel et al., 2007). This specific gas phase can be traced by the rotational H2 transitions seen in the mid-infrared (MIR); (iii) The hot molecular phase, K, mainly traced by the roto-vibrational H2 transitions in the near-infrared. This gas phase represents only of the mass of the cold molecular gas phase (Dale et al., 2005; Emonts et al., 2014; Pereira-Santaella et al., 2016). Spatially resolved observations of the structure and kinematics of both cold and hot molecular gas emissions in the central regions of galaxies have been meticulously obtained through ALMA and near-infrared integral field unit observations (Lamperti et al., 2022; Ramos Almeida et al., 2022; Speranza et al., 2022; Bianchin et al., 2022; Riffel, 2021; Riffel et al., 2023b).
@@ 3 Results and discussions
The emission of the outflow component (bottom panels of Fig. 2) is spatially resolved (with FWHM of its flux distribution of about 2.4 times that of the PSF) and concentrated mostly around the MIR nucleus and radio core. In addition, we simulated a rotating disk convolved with the PSF to assess the impact of beam smearing on the observed line profiles, which was found to be negligible (refer to the Appendix A). This component is blue-shifted relative to the systemic velocity of the galaxy and presents values of up to 600 km s-1. The median velocity and values are 70 km s-1 and 455 km s-1 for the [Ar ii] and 50 km s-1 and 430 km s-1 for the H2, respectively. From the observed H2 emission and kinematics, we derive the warm molecular gas outflow properties in UGC 8782.
We measure the gas mass and temperature using the transitions H2 (0-0) S(1),H2 (0-0) S(2) , H2 (0-0) S(3), H2 (0-0) S(4) and H2 (0-0) S(5), as our data does not include the higher order transitions, as well as the H2 (0-0) S(0) transition. We correct the observed fluxes for extinction using a visual extinction of mag, the mean value for the outflow component (Riffel et al., 2023a) and adopting the extinction law G23 (Gordon et al., 2023, 2009; Fitzpatrick et al., 2019; Gordon et al., 2021; Decleir et al., 2022).
@@
Given the AGN bolometric luminosity of erg s-1 (Riffel et al., 2023a), the molecular outflow demonstrates maximum kinetic coupling efficiencies () ranging from 2–5% of the AGN luminosity, the kinectic power of the warm and hot phase are respectively erg s-1 and erg s-1. This range far surpasses the minimum required for effective AGN feedback in suppressing star formation (Hopkins & Elvis, 2010), particularly when considering the potential presence of a cold molecular gas reservoir that may also contribute to the outflow. Nevertheless, the outflows in UGC 8782 are likely mechanically driven by the radio jet, as the kinetic power of the outflow is one order of magnitude lower than the object’s radio source power of 1043 erg s-1 (Lanz et al., 2015). Considering a constant outflow rate the warm and hot H2 gas reservoir would be depleted in around Myr.
@@ 4 Conclusions
We reported the observations of warm molecular gas spatially resolved outflows from the AGN of UGC 8782. We find that the mass outflow rate in warm molecular gas is substantially higher than that observed in ionized gas across all radii, with coupling efficiencies exceeding the minimum required for AGN feedback to be effective in quenching star formation. The observed molecular outflows in UGC 8782 of up to 6 M⊙yr-1 are able of redistributing all the warm molecular gas available in the inner 2 kpc of the UGC 8782 in only 1 Myr.